Venus, with an orbit of radius 1.08 x 1011 m, takes 225 days to orbit the sun. What is the average speed of
Venus in metres per second? Determine the magnitude of the average velocity after it has completed half a
revolution around the sun. Also, determine the magnitude of the average acceleration after it has completed a
quarter revolution around the sun.
1
Expert's answer
2019-09-23T09:14:40-0400

1.


"V=\frac{2\pi R}{T}=\frac{2\pi (1.08\cdot 10^{11})}{225\cdot 24\cdot 3600}=34900\frac{m}{s}"

2.


"V_a=\frac{2 R}{0.5T}=\frac{4 (1.08\cdot 10^{11})}{225\cdot 24\cdot 3600}=22200\frac{m}{s}"

3.


"a=\frac{(V)^2}{R}=\frac{(34900)^2}{1.08\cdot 10^{11}}=0.0113\frac{m}{s^2}"